OpenStudy (anonymous):

Between which two integers does √12 lie? A. 3 and 4 B. 5 and 7 C. 11 and 13 D. 143 and 145

OpenStudy (freckles):

hint: 16>12>9

OpenStudy (freckles):

now take the square root

OpenStudy (anonymous):

sorry i dont get it D:

OpenStudy (anonymous):

\[\sqrt{12}=3.4641 \]A.

OpenStudy (anonymous):

can you walk me through it?

OpenStudy (anonymous):

is it A?

OpenStudy (freckles):

\[\sqrt{16}>\sqrt{12}>\sqrt{9}\]

OpenStudy (anonymous):

Thankyou to both for helping me :D x I get it now.

OpenStudy (freckles):

you could square everything instead of that question ask this: Between which two integers does 12 lie? A) 9 and 16 B) 25 and 49 C) 121 and 169 D) 143^2 and 145^2 (lol couldn't do these in my head)

OpenStudy (anonymous):

oh hahaha, thankyou btw c: